# Privacy Technologies for Financial Intelligence

### v1

@@ Abstract
Financial crimes like terrorism financing and money laundering can have real impacts on society, including the abuse and mismanagement of public funds, increase in societal problems such as drug trafficking and illicit gambling with attendant economic costs, and loss of innocent lives in the case of terrorism activities. Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle is usually distributed across a network of financial institutions, regulators, and law-enforcement agencies and they cannot be easily shared due to privacy constraints. Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology. This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection.
@@ 1 Introduction
Our economy depends on a well functioning banking and financial system that has the trust of the general population. However in recent years, the reputation of some major banks has been severely damaged as criminal elements have exploited systemic weaknesses to launder huge sums of money, including transfers to terrorist organisations [Lyn22, LF23]. Besides reputational damage to financial institutions, complex financial crimes like terrorism financing and money laundering can have real direct impacts on society too, including
  • abuse and mismanagement of public funds;
  • increase in societal problems like drug trafficking, illicit trade and gambling, people smuggling, with attendant economic costs; and
  • loss of innocent lives in the case of terrorism activities.
Complex financial crimes can be hard to detect primarily because data related to different pieces of the overall puzzle are usually distributed across a network of financial institutions, regulators, and law-enforcement agencies [LCSH19]. The problem is also rapidly increasing in complexity because of new emerging value-transfer platforms like crypto-currencies and real-time settlement systems.
In early 2016, AUSTRAC, Australia’s Financial Intelligence Agency, set up the Fintel Alliance [^1] https://www.austrac.gov.au/partners/fintel-alliance, an organisation that includes AUSTRAC and the major financial institutions and law-enforcement agencies, to improve public-private collaboration in combating financial crimes. Since then, financial intelligence agencies around the world have started building public-private data-sharing platforms to share and disrupt criminal entities engaging in financial crimes, including the Joint Money Laundering Intelligence Task Force (JMLIT) [^2] https://www.nationalcrimeagency.gov.uk/what-we-do/national-economic-crime-centre in UK, and the Financial Crime Enforcement Network Exchange [^3] https://www.fincen.gov/resources/financial-crime-enforcement-network-exchange in the United States. These public-private platforms are highly effective against known threats.
@@
However, their discovery functions to find as-yet unknown risks in the financial system are still not as strong as they can be, primarily because of privacy constraints around large-scale data matching and analytics between organisations. Recent advances in Privacy-Preserving Data Matching and Machine Learning provide an opportunity for regulators and the financial industry to come together to solve the risk-discovery problem with technology. This paper provides a survey of the financial intelligence landscape and where opportunities lie for privacy technologies to improve the state-of-the-art in financial-crime detection. Our paper complements the material covered in [Max20], which explored global case studies of public and private sectors employing cryptographic-based privacy technologies to foster collaborative financial intelligence. [^4] https://www.future-fis.com/the-pet-project.html
@@ 2 Financial Intelligence
Financial intelligence is a highly specialised field that is not well understood by many. We provide an introduction to the financial intelligence ecosystem in this section, including a description of the current state, some existing limitations and opportunities.
@@ 2.1 The Financial Intelligence Ecosystem
Financial Intelligence Units (FIUs) around the world are governed by anti-money laundering and counter-terrorism financing (AML/CTF) regimes that are quite similar in nature [Mad11, Gil04]. Generally speaking, FIUs collect data like threshold transaction reports, international fund transfer instructions, and suspicious matter/activity reports from Reporting Entities (REs), which include banks, money remitters, casinos, law firms, real-estate companies, and financial companies. (In a country the size of Australia, there are about 10,000 REs in the financial system.) FIUs may also get data about entities of interest from partner agencies like law-enforcement agencies and partner FIUs in other countries. Using all these collected data, analysts in FIUs produce intelligence products that either report on specific threats that can be actioned by law enforcement, or describe macro threats and mitigation strategies that financial institutions should adopt in their internal controls. Figure 1 shows the general information flow from reporting entities to FIUs and, from there, to law-enforcement agencies.
Figure 1: A reporting entity is a company that provides financial services that are regulated under AML/CTF laws; KYC = Know Your Customer; ECDD = Enhanced Customer Due Diligence; TTRs = Threshold Transaction Reports; IFTIs = International Fund Transfer Instructions; SMRs = Suspicious Matter Reports; API = Application Programming Interface.
There are safeguards and risk monitoring all over the ecosystem, inside each reporting entity and within the FIU, and there is a reasonable mix of automation and human involvement in those risk-monitoring and risk-reduction processes. The parts of the diagram that are painted in brown are areas of the financial intelligence ecosystem where advanced analytics like predictive and prescriptive modelling can play an important role, mostly because the data volume that need to be processed at those places are enormous and the relationships that need to be mastered and managed, including cause and impact relationships, are often complex.
@@
Generally speaking, all the data received by an FIU are pushed through an entity resolution step to identify and merge records that are highly likely to refer to the same entity [ZNCC18]. This is usually a highly non-trivial computational problem because of the sheer number of parties involved in financial transactions (billions for a typical FIU), to arrive at a table like Figure 2, where the rows are entities (either individuals, companies, or groups of related individuals/companies) and the table show their behaviour in the financial system as seen through the lens of different reporting entities. In some cases, we also have explicit risk indicators for these entities. Given such a table, an FIU equipped with data science expertise and big data analytics platforms can use profiling and statistical machine learning techniques, both supervised and unsupervised at different levels of granularity from individual transactions to single entities and networks of entities, to uncover risks in the financial system and act on them. Examples of such work include [SWC+16, SWZ+17, SZC+17].
@@
Figure 2: The entity-resolved data and risk indicators available to a typical FIU.
@@
A key thing to understand about financial intelligence is that financial crimes cannot really be detected from financial transaction data alone. Here, in a nutshell, are the sort of things we can infer about anomalies in financial transactions [Mad11, RT04]:
@@
    the source of fund cannot be explained;
@@
    there are rapid movements of funds through different channels, which may indicate intention to hide the money trail;
@@
    the financial transaction behaviour is inconsistent with a customer’s expected business profile.
These sorts of activities are not illegal on their own, of course. It is only when these activities accompany a predicate crime — tax evasion, welfare fraud, drug trafficking, etc — that criminality can be established and treated. For that reason, one needs law-enforcement presence throughout the ecosystem to supply leads and the risk contexts for the whole system to work well. In particular, pure discovery work in financial transaction data is quite difficult but commercial entities of interest databases like LSEG World-Check and external datasets like company registration data can be useful in adding context to financial data.
That is the current state. The overall system is far from perfect but, by and large, the regime works reasonably well in deterring, detecting, and disrupting financial crimes. The main limitation is in the links that are not there in the ecosystem. First of all, a lot of financial transaction data, including domestic interbank transfers regardless of amount, are not reported by default to an FIU. Secondly, there are literally tens of thousands of reporting entities in a mature economy like Australia and these reporting entities are largely limited in what intelligence and data they can share with each other because of legal impediments around sharing SMRs between reporting entities to minimise the risk of (accidental) tipping off. There is also not a lot of feedback loops built into the system, including that between reporting entities and an FIU, and that between an FIU and law-enforcement agencies.
The important limitations in the current state can thus be mostly traced to the following two issues:
  • Complex financial crimes typically involve transactions that involve multiple financial institutions in multiple geographies, going through multiple payment channels, some of which are unregulated.
  • FIUs and reporting entities are all working with data silos and no one has a complete view of activities across the financial system.
To appreciate the extent of the problem, one only needs to look at the difficulty of trying to answer a simple question like finding all entities that exhibit a certain cross-border transaction pattern like that shown in Figure 3.
Figure 3: A simple hard problem
The question, and many more complicated others we would like to ask, simply cannot be answered today because the data needed to answer them sit in different organisations in different countries, governed by different legislation.
2.2 Privacy Technologies As A Way Forward
So what can we do? Figure 4 shows a possible way forward using distributed confidential computing technologies, assuming FIUs and partner agencies are willing to work together but without needing to make wholesale changes to AML/CTF legislations everywhere.
@@ 2.2 Privacy Technologies As A Way Forward
The obvious obstacle to different FIUs and partner agencies sharing large volumes of data with each other is privacy concerns. In the proposed scheme, instead of entity resolution, we have all the participating agencies contributing data that are first matched using privacy-preserving record linkage (PPRL) algorithms [CRS20]. In a PPRL algorithm, the records in different databases that represent or belong to the same entity are identified in a privacy preserving manner without revealing any plaintext values to the participating parties. The matched data are then encrypted using a homomorphic encryption scheme [Hal17] (as described in Section 3.2), which allow mathematical operations to be performed on encrypted data, to arrive at a table similar to what we showed earlier, with two distinctions:
  • The values in the table are all encrypted and the columns of the table are stored in a distributed way across different databases in different agencies.
  • The point of using homomorphic encryption schemes is that it is still possible to run a range of risk-modelling algorithms over the encrypted data, and once we get to a small set of entities of interest, decrypting those entities’ data can be done without raising any privacy concerns for the general public.
Having the data distributed in different data sources does complicate the running of risk-modelling algorithms somewhat but the added security and scalability make it a worthwhile trade-off to make.
There are basically three fundamental computational problems to solve, in increasing order of difficulty:
@@
    How can two FIUs, each of which has a list of suspicious entities, share only the common suspicious entities with each other, without sharing their entire list with each other? More generally, how can such sharing be done between multiple FIUs?
@@
    Given a criminal typology that can be described as a graph pattern, where nodes denote entities and edges denote relationships, how can we detect all entities that satisfy the criminal typology when the underlying graph span multiple financial institutions?
@@
    How can we perform model learning, either supervised or unsupervised, over a set of entities whose financial data and behaviour are spread across multiple financial institutions?
In the rest of this paper, we will look at how aspects of that future state can be realised through different privacy technologies.
@@ 3 Privacy Technologies
The purpose of this section is to briefly introduce the most common privacy technologies that are applicable to financial intelligence. These techniques are based on different mathematical foundations to serve different data privacy and security purposes. Secure multiparty computation, through the use of secret sharing schemes, permits multiple parties to jointly compute a function without divulging each party’s secret information to the others. Homomorphic encryption tackles this issue from a different perspective, by encrypting sensitive data in a way that allows arithmetic operations to be performed directly on the ciphertexts. Some privacy issues could arise through reverse-engineering the algorithm or query outcomes [DN10, KM11]. To address this, the differential privacy framework can be used to provide individuals with guaranteed plausible deniability by adding suitably calibrated noise to query outcomes. We will also describe federated learning, which is needed for multiple parties to jointly learn a machine learning model from distributed data. We will start the section with hashing, a foundational technique for many privacy technologies.

### v2

@@ Abstract
Financial crimes like money laundering and terrorism financing can have significant impacts on society, including loss of trust in the integrity of the financial system, misuse and mismanagement of public funds, increase in societal problems like drug trafficking and illicit gambling, and loss of innocent lives due to terrorism activities. Effective detection of complex financial crimes remains a formidable challenge for regulators and financial institutions because the critical data needed to establish patterns and criminality are often dispersed across multiple organisations and cannot be linked due to privacy constraints around large-scale data matching. Recent advances in privacy and confidential computing technologies, which enable private and secure data analysis across organisations, offer a promising opportunity for regulators and the financial industry to come together to enhance their collaborative risk detection while maintaining privacy standards. This paper, through a survey of the financial intelligence ecosystem, seeks to identify opportunities for the utilisation of privacy technologies to improve the state-of-the-art in financial-crime detection.
keywords:
Financial intelligence , Anti-money laundering , Privacy technology
[…]
[inst1]organization=School of Information Technology, Deakin University, state=VIC, country=Australia
[…]
[inst2]organization=Data61, CSIRO, state=ACT, country=Australia
[…]
[inst3]organization=School of Computing, Australian National University, state=ACT, country=Australia
@@ 1 Introduction
A robust and trustworthy banking and financial system is essential for the stability and proper functioning of any mature economy. However, in recent years, the reputation and integrity of several major financial institutions have been seriously undermined as organised criminal networks have exploited loopholes and systemic vulnerabilities to launder vast sums of money, some of which have even been linked to terrorist financing [39, 37]. Beyond the reputational harm inflicted on financial institutions, sophisticated financial crimes like money laundering and terrorism financing have profound societal consequences, including:
  • 1.
    mismanagement and misuse of public funds;
  • 2.
    financial losses for institutions due to increased compliance costs and regulatory penalties;
  • 3.
    escalation of societal issues such as drug trafficking, human smuggling, and illegal gambling, which carry significant economic burdens;
  • 4.
    the tragic loss of innocent lives in the case of terrorist activities.
Globally important cases like Binance’s \$4 billion fine [^1] Binance’s \$4 billion fine for failing to maintain an effective anti-money-laundering (AML) program and other violations against the US Bank Secrecy Act, UBS’s \$4.2 billion fine by Swiss, US and French regulators for facilitating tax evasion and inadequate AML controls, and Goldman Sachs’s \$2.9 billion fine [^2] Goldman Sachs’s \$2.9 billion fine for the 1MDB money-laundering and bribery violations, have certainly brought these issues into the public consciousness and discourse. In Australia, the government’s successful court proceedings against the Commonwealth Bank of Australia, Westpac Banking Corporation, and Crown Casinos for serious breaches of the AML/CTF Act, and their ongoing legal actions against the operators of the Star Casinos and online betting sites like Ladbrokes and Neds highlight the continued challenges faced by regulators and financial intelligence agencies in combating financial crimes. [^3] Lists of enforcement actions taken
In response to these challenges, AUSTRAC, Australia’s Financial Intelligence Agency, established the Fintel Alliance [^4] Fintel Alliance in 2016 to foster collaboration between government agencies, financial institutions, and law enforcement. A number of similar initiatives have also emerged globally, including the UK’s Joint Money Laundering Intelligence Task Force [^5] UK’s Joint Money Laundering Intelligence Task Force and the U.S. Financial Crime Enforcement Network Exchange [^6] U.S. Financial Crime Enforcement Network Exchange. These partnerships have proven effective in addressing known threats, but they still face formidable challenges in detecting complex financial crimes because the critical data needed to establish patterns and criminality are often dispersed across multiple organisations and cannot be linked due to privacy constraints around large-scale data matching [36]. The problem is further compounded by the growing complexity of emerging value-transfer mechanisms such as cryptocurrencies and real-time payment systems.
@@
Recent innovations in Privacy-Preserving Data Matching and Machine Learning, which enable private and secure data analysis across organisations, offer a promising opportunity for regulators and the financial industry to come together to enhance their collaborative risk detection while maintaining privacy standards. This paper, through a survey of the financial intelligence ecosystem, seeks to identify opportunities for the utilisation of privacy technologies to improve the state-of-the-art in financial-crime detection. Our paper complements the material covered in [41], which explored global case studies of public and private sectors employing privacy technologies to foster collaborative financial intelligence.
@@ 2 Financial Intelligence
Financial intelligence is a highly specialised field that is not well understood by many. This section provides an introduction to the financial intelligence ecosystem, including a description of the current state, some existing limitations and opportunities.
@@ 2.1 The Financial Intelligence Ecosystem
Financial Intelligence Units (FIUs) around the world are governed by anti-money laundering and counter-terrorism financing (AML/CTF) regimes that are quite similar in nature [40, 28]. Reporting Entities (REs) are companies that provide designated financial services that are regulated under AML/CTF laws and typically include banks, money remitters, digital currency exchanges, casinos, and increasingly law firms and real-estate companies. (In a country the size of Australia, there are about 10,000 REs in the financial system.) These REs are required by law to maintain know-your-customer (KYC) and anti-money laundering programs to make sure they can verify their customers’ identities and prevent their designated services from being used for illicit purposes. Certain types of customer transaction data, including threshold transaction reports (TTRs) for cash transactions exceeding \$10k and all international fund transfer instructions (IFTIs) regardless of amount, are automatically sent to the relevant FIUs on a regular basis. In addition, REs can also submit suspicious matter reports (SMRs) or suspicious activity reports (SARs) to FIUs if they detect suspicious customer activities through their internal controls. Adding to all these data from REs, FIUs may also get reports about specific individuals or entities of interest from law-enforcement agencies and partner FIUs in other countries. Using all these collected data, analysts in FIUs produce intelligence products that either report on specific threats that can be actioned by law enforcement agents, or describe macro threats and mitigation strategies that financial institutions should adopt in their internal controls. Figure 1 shows the information flow from reporting entities to FIUs and, from there, to law-enforcement agencies and the limited feedback loops in the system.
Figure 1: The figure shows a simplified schematic view of a financial intelligence ecosystem.
As we can see, there are risk-monitoring and risk-reduction mechanisms built into many different parts of the financial intelligence ecosystem, both within reporting entities in a distributed way and more centrally within the FIU. Many of these organisations also have large-scale automated transaction-monitoring systems that are augmented by manual analyst investigations and formal governance oversight. The parts of Figure 1 coloured in brown represent areas of the financial intelligence ecosystem where artificial intelligence and machine learning can have a significant impact. This is primarily because the sheer quantity of data that must be processed in these locations is immense, and the connections and relationships that need to be distilled and understood, including causal and deliberately hidden linkages, are often complex and intricate.
@@
Broadly speaking, all data received by an FIU are typically first processed through an entity-resolution step to identify and consolidate records that are highly probable to relate to the same underlying entity [65]. The aim is to produce a dataset like that shown in Figure 2, where each row represents an entity (either persons, corporations, or clusters of related individuals/companies) and the columns illustrate their activities within the financial ecosystem as observed through the lens of the financial institutions whose services they engage with. A small set of such entities would also have explicit risk indicators flagged against them based on, for example, suspicious matter reports, sanctions list and other criminal intelligence data. Given a dataset like Figure 2, an FIU equipped with appropriate technical expertise and advanced data-analytics platforms can employ profiling and statistical machine learning methodologies, both supervised and unsupervised, at varying levels of granularity from individual transactions to single entities and networks of entities, to try to uncover potential risks in the financial system and take appropriate action. Examples of such work include [55, 56, 57].
@@
Figure 2: The entity-resolved data and risk indicators available to a typical financial intelligence agency.
@@
A crucial aspect to understand about financial intelligence is that financial crimes cannot be definitively identified solely from financial transaction data. Here, at a high level, are the types of anomalies we can infer from financial transaction data [40, 52]:
@@
    the origin of funds cannot be explained or is obscure;
@@
    there are rapid transfers of money through various channels, potentially indicating an attempt to obscure the money trail;
@@
    the financial transaction patterns are inconsistent with a customer’s anticipated personal or business profile.
It is important to note that these activities, in isolation, are not inherently illegal. Criminality can only be established and addressed when these activities are linked to a predicate crime, such as tax evasion, welfare-payment fraud, drug trafficking, or similar crimes. To ensure the entire system functions effectively, law enforcement presence is essential throughout the financial intelligence ecosystem to provide leads and supply appropriate risk contexts. Specifically, pure discovery work within financial transaction data presents considerable challenges, but commercial databases of entities of interest, such as LSEG World-Check, and external datasets like company registration information can prove valuable in adding context to financial data.
That is the current state. The overall system is not perfect by any means. Nevertheless the regime is reasonably effective in deterring, detecting, and disrupting financial crimes, as demonstrated by the many successful prosecutions listed in Section 1. The main limitation of the current state is in the links that are there in the physical financial system but remain missing in the financial intelligence ecosystem. To begin with, a lot of financial transaction data, including domestic interbank transfers and crypto currency transactions, are not reported by default to an FIU and that is a major blind spot for analysts. Compounding the issue, the many thousands of reporting entities in a mature economy like Australia are heavily handicapped in their ability to share intelligence and data with each other because of legal impediments and risks around accidental tip-offs. One can also argue that there is simply not enough feedback loops built into the system, including that between reporting entities and an FIU, and that between an FIU and its partner law-enforcement agencies and foreign FIUs.
In summary, the key limitations in the current state of financial intelligence can be traced to the following issues:
  • 1.
    Sophisticated and complex financial crimes typically involve transactions that span multiple financial institutions in multiple geographies, going through multiple payment channels, some of which remain unregulated.
  • 2.
    Criminality cannot be established through analysis of financial transaction data alone; actual links to predicate crimes are necessary.
  • 3.
    FIUs, reporting entities, and law-enforcement agencies are all mostly working within data silos and it is hard for any one organisation to get a complete view of financial and criminal activities across the entire system.
2.2 A Way Forward using Privacy Technologies
To overcome the key limitations of the current state of financial intelligence, Figure 3 outlines a potential solution that leverages distributed confidential computing technologies. The assumption is that FIUs, reporting entities and law-enforcement agencies are willing to collaborate in good faith without requiring sweeping changes to AML/CTF legislation globally.
@@ 2.2 A Way Forward using Privacy Technologies
In the proposed approach, rather than traditional entity resolution, data from all participating organisations are first matched using privacy-preserving record linkage (PPRL) algorithms [18]. These algorithms identify records across different databases that correspond to the same entity in a manner that safeguards privacy, ensuring no plaintext values are disclosed to any party involved.
Once the records are matched, the data are encrypted using either homomorphic encryption (Section 3.2) or secure multiparty computation schemes (Section 3.1). These encryption methods allow mathematical operations to be performed directly on encrypted data, resulting in a table similar to those discussed earlier but with two key differences:
  • 1.
    All values in the table are encrypted, and the columns are distributed across multiple databases managed by different agencies.
  • 2.
    AI and machine learning algorithms for detecting risks and suspicious entities can still be run on the encrypted data using homomorphic encryption or secure multiparty computation. When a small set of entities of interest is identified, their data can be decrypted without compromising the privacy of the broader population.
Distributing encrypted data across multiple organisations and databases does introduce some complexity to AML/CTF risk-modelling workflows, but the approach offers a pathway for collaboration while addressing privacy concerns and ensuring compliance with existing legal frameworks.
To help us navigate the opportunities offered by privacy technologies for financial intelligence, it is essential to understand that there are three fundamental computational problems to address in AML/CTF risk modelling, in increasing order of technical complexity:
@@
    Private Intersection Identification: How can two FIUs, each maintaining its own list of suspicious entities, securely exchange overlapping data points without exposing their full datasets with each other? How might this process scale across multiple FIUs?
@@
    Cross-Institutional Pattern Detection: Given a criminal typology defined as a graph pattern (where nodes represent entities and edges denote relationships), how can we identify all entities matching this pattern when the underlying graph pattern spans multiple financial institutions?
@@
    Federated Model Learning: How can supervised or unsupervised machine learning models be trained on behavioural and transactional data distributed across multiple financial institutions, where no single entity holds complete datasets?
These three computational problems are distilled from the authors’ experience working both directly and indirectly with Australia’s financial intelligence and banking communities over an extended period of time, and backed up by literature review.
In the rest of this paper, we will first survey some key privacy technologies in § 3 before providing in § 4 illustrative examples and detailed pointers to how such privacy technologies can be used to address the three computational problems listed above. Solutions to these three computational problems can be, and are often, used as building blocks to develop more complete systems to address challenging real-world financial intelligence problems, some of which we survey in § 5.
@@ 3 Privacy Technologies
In the above, we highlighted a central challenge in multiparty financial intelligence, how can multiple parties work together effectively when none of them can be assumed to be fully trusted? We also noted that privacy technologies can serve as a technological means to enable such collaboration, allowing multiple parties to jointly combat financial crime under specific cooperation models. We refer to this form of multiparty financial intelligence conducted under explicit data privacy and security constraints as privacy-preserving financial intelligence.
Privacy-preserving financial intelligence should be understood as a system level concept rather than a single method. Different privacy technologies play different roles within this system, but they share a common objective of enabling analysis and modelling, while maintaining data security and privacy guarantees.
This section begins with secure multiparty computation (MPC), as it is one of the most widely adopted mechanisms for protecting data in collaborative settings. MPC enables multiple parties to jointly perform computations without revealing their private inputs. Building on this foundation, we introduce homomorphic encryption (HE), which supports MPC by enabling computation over encrypted data and is therefore often used as a cryptographic primitive in MPC designs. We then move beyond MPC centric settings and discuss several problem scenarios where MPC is not the primary solution. For these cases, we introduce privacy technologies that are widely used in practice, namely hashing, federated learning (FL), and differential privacy (DP), each addressing privacy risks from a different perspective.
3.1 Secure Multiparty Computation
The goal of secure multiparty computation (MPC) is to enable multiple parties to jointly evaluate a pre-agreed function without revealing their individual private inputs. At its core, MPC removes the need to trust the computation process itself. Under an MPC protocol, each participant learns only the final output (or an authorised share of it) and cannot infer additional information about other parties’ inputs beyond what is inherently revealed by that output. In doing so, MPC shifts cross-institutional collaboration away from data centralisation or reliance on a trusted third party, and instead enforces privacy and correctness directly at the protocol level through cryptographic mechanisms.
More generally, MPC can be viewed as a general-purpose framework for privacy-preserving computation. Whenever a collaborative task can be formalised as a function, MPC provides a mechanism for jointly evaluating that function without requiring participants to share their raw data.
Because MPC is a strict security framework, its practical deployment depends on clearly defined security guarantees. In particular, privacy ensures that an adversary learns nothing about an honest party’s input beyond what follows from the output itself. Correctness guarantees that the protocol produces the correct result, even in the presence of adversarial behaviour. Stronger notions may also be required in certain settings: guaranteed output (or robustness) ensures that adversaries cannot prevent honest parties from obtaining the result by aborting the protocol, while fairness requires that an adversary receives the output if and only if the honest parties do.
In addition to these properties, MPC protocols are analysed under explicit adversarial models. In the semi-honest (honest-but-curious) model, adversaries follow the protocol specification but attempt to extract additional information from the execution. In the malicious model, adversaries may deviate arbitrarily from the protocol, for example by sending malformed messages, aborting prematurely, or manipulating inputs and outputs. The assumed adversarial model directly influences both the strength of the security guarantees and the complexity of the resulting protocol.
In the context of financial intelligence, these capabilities can be illustrated through a common investigative scenario (discussed further in Section 4.1). Different financial institutions (FIs) maintain their own customer lists, while a financial intelligence unit (FIU) may need to determine which entities on its watchlist have relationships with which banks. Here, the FIU and the banks act as participating parties, their private inputs are their respective customer lists, and the agreed computation is the intersection between the FIU’s list and those of the banks. MPC enables this joint analysis to be conducted without exposing any party’s complete customer list.
3.2 Homomorphic Encryption
Within the MPC framework, privacy protection is not achieved by simply hiding the output, but by ensuring that computation itself can be performed without exposing sensitive inputs. For this reason, MPC protocols typically rely on multiple cryptographic mechanisms rather than a single technique. Homomorphic encryption (HE) is one such foundational capability.
The defining feature of HE is that it allows certain computations to be performed directly on encrypted data, producing results that, once decrypted, are equivalent to those obtained from computations on plaintext. In MPC settings, this property enables parts of a computation to be executed without revealing intermediate values, thereby reducing direct data exposure between participants.
$\text{Dec}(\text{secret key},\text{Enc}(\text{public key},m_{1})\circ\text{Enc}(\text{public key},m_{2}))=m_{1}\diamond m_{2},$ (1)
From this perspective, HE is not a replacement for MPC, but rather one of its supporting building blocks. In practice, HE is often combined with other mechanisms, such as secret sharing or oblivious transfer, to securely encapsulate specific computation steps while balancing privacy guarantees, communication overhead, and system complexity. At the same time, HE can also be used independently of a full MPC framework. In scenarios where the computation structure is relatively simple, for example, when one party holds the data and another performs the computation, HE allows encrypted data to be processed externally without revealing the underlying inputs. This enables analysis or evaluation without requiring raw data sharing.
HE schemes are commonly classified according to the types of operations they support and the number of operations that can be performed before decryption becomes unreliable. The most general form is Fully Homomorphic Encryption (FHE), which supports an unbounded number of operations on encrypted data. The practical feasibility of FHE was established following Gentry’s breakthrough construction [27], which sparked significant research interest, particularly in lattice-based cryptography, a field now widely regarded as central to quantum-resistant encryption [2].
It is also important to note that HE schemes are inherently probabilistic, which means encrypting the same plaintext multiple times produces different ciphertexts. This property ensures semantic security, which prevents adversaries from inferring information about the plaintext from ciphertext patterns.
In financial intelligence applications, such standalone use of HE may arise in outsourced computation of sensitive transaction features, risk metrics, or aggregate statistics. Financial institutions can retain full control over their data while leveraging external compute resources or cross-organisational collaboration, without disclosing plaintext information.
Despite its strong privacy guarantees, the applicability of HE remains constrained by performance considerations, supported computation types, and overall system complexity. As such, HE is best understood either as a specialised independent capability within the privacy-preserving toolkit, or as a key component within more complex MPC-based solutions.